OpenStudy (alexmc8):

Find the exact values of sin A and cos A. Write fractions in lowest terms. A right triangle ABC is shown. Leg AC has length 36, leg BC has length 48, and hypotenuse AB has length 60. sin A = three fifths; cos A = four fifths sin A = Afour thirds; cos A = three fourths sin A = five fourths; cos A = five thirds sin A = four fifths; cos A = three fifths

OpenStudy (alexmc8):

Please help. I will give a medal

OpenStudy (eliesaab):

This a right triangle in C. Sin A=36/60=3/5 How much is the Cos A

OpenStudy (eliesaab):

Sin =(opposite side)/Hypotenuse Cos=(adjacent side)/Hypotenuse

OpenStudy (eliesaab):

It is a right triangle since \[ 60^2=48^2+36^2=3600 \]

OpenStudy (alexmc8):

cos A = four fifths

OpenStudy (eliesaab):

Yes

OpenStudy (alexmc8):

Thank you for helping (: